What is the boiling point of a solution that contains 1.25 mol CaCl2 in 1400 g of water
1
Expert's answer
2020-04-20T15:08:38-0400

Law of solutions (Raul's) dTboillingpoint=E "\mu"

The molality is n(CaCl2)/m(H2O)=0.892 mole/kg

dT=0.52*0.892=0.464 C